# Observational Probes of the Neutron Star Equation of State with Hyperons, Bosonic Dark Matter, and Quark Matter

### v2

@@ 1 Introduction
In our prior work, Shahrbaf et al. (2022), we established the lower-mass limit for S ensuring that the possible emergence of S as a DM candidate remains consistent with observational constraints derived from NSs. Our findings disclosed the feasibility of S appearance in the hadronic phase of NS, demonstrating that a minimum mass of $m_{\texttt{S}}=1885$ MeV fulfills the necessary conditions that S does not appear at sub-saturation densities. In the present work, we adopt the same approach to ascertain the upper limit for the mass of S, assuming that the interaction strength with BM is fixed at its lowest value. This provides a reliable assumption about the DM-ordinary matter interaction. We found that $m_{\texttt{S}}=2040$ MeV represents the upper mass limit for S particles, ensuring that the observable characteristics of the DM admixed NS are in favor of their astrophysical boundaries.
@@
Furthermore, we perform a Bayesian analysis (BA) to systematically determine the posterior probability distribution of the allowed mass range of S, accounting for both observational uncertainties and model assumptions. By applying multiple constraints, the Bayesian framework provides more precise limits on the allowed bosonic DM parameter space within the context of NS physics. For our analysis, we consider fundamental constraints on NSs, mass, radius, and tidal deformability along with two extreme cases: HESS J1731-347 (Doroshenko et al., 2022), a light and ultra-compact NS, and PSR J0952–0607 (Romani et al., 2022), the most massive confirmed NS, known as the “Black Widow” (BW) pulsar. The constraints are first applied independently to infer the respective posterior distributions of the S mass. Subsequently, these are combined to yield the final posterior distribution.
@@
The paper is organized as follows: In Section 2, we introduce our model to describe the hadronic phase with hyperons and DM as well as the QM phase. Our approach for combining these two phases and construct a crossover phase transition is also discussed in this section. Our findings and in-depth discussions concerning the thermodynamic properties of the resulting hybrid stars are presented in Section 3. In Section 4, the impact of NSs constraints to find the upper limit of the mass of DM particles has been investigated. Our BA approach based on the NS observations is presented in Section 5. Finally, we summarize our work and draw conclusions in Section 6.
@@ 2.1 Hadronic Matter Equation of State Including Hyperons and Dark Matter
In the DD2Y-T model, all particles are treated as quasi-particles with effective mass and effective chemical potential. The interactions in this model exclusively entail meson exchange interactions, and the couplings to $\sigma$, $\omega$, $\rho$, and $\phi$ mesons depend on the density of the medium. The density-dependent couplings are accurately adjusted to faithfully reproduce the characteristics of atomic nuclei, as established by earlier works Typel & Wolter (1999); Typel (2005); Typel et al. (2010).
The DD2Y-T model has been employed to assess the properties of NSs in both high-temperature, or ”hot”, and low-temperature, or ”cold”, hyperonic configurations (Stone et al., 2021). Notably, it is essential to highlight that this model successfully addresses the ’hyperon puzzle’, a longstanding problem that has been extensively deliberated in the literature (Sedrakian et al., 2020; Choi et al., 2023; Tolos, 2025; Fujimoto et al., 2024; Tong et al., 2025). It is considered one of the most challenging issues within the domain of hypernuclear physics, and various works have tried to address the solutions to this puzzle (Maslov et al., 2015; Shahrbaf & Moshfegh, 2019; Shahrbaf et al., 2020a, b; Li & Sedrakian, 2023; Tsiopelas et al., 2024; Jangal et al., 2025).
@@
As established in Shahrbaf et al. (2022), we have demonstrated that a constant mass assumption for the S particle is incompatible with the structural stability of NS. Consequently, we introduced a linear density-dependent mass shift for the S particle, with a slope parameter harmonized with the corresponding mass shifts observed in other baryons. Detailed constraints regarding the lower threshold of the mass and the slope parameter have been thoughtfully presented and illustrated in Figure 8 of Shahrbaf et al. (2022). In our approach
$S_{S}=-\Delta m_{S},$ (1)
@@
$\Delta m_{S}=m_{S}x_{S}\frac{n_{B}}{n_{0}}.$ (3)
where $n_{B}$ is the baryonic density, $x_{S}$ is a slope parameter and $n_{0}=0.15$  fm-3 is the nuclear saturation density. A finite density of S exclusively emerges at zero temperature when the condition $m^{*}_{S}=\mu^{*}_{S}$ is met, leading to the occurrence of Bose-Einstein condensation. The scalar density for S then is equal to the vector density and reads
@@
where $\mu_{S}^{\ast}=\mu_{S}-V_{S}$ is the effective chemical potential and $\mu_{S}=2\mu_{B}$ is the chemical potential of S particle. $V_{S}$ is the vector potential for the S particle and reads
@@ 2.3 Hybrid Equation of State
In the context of constructing a phase transition to QM, traditional Maxwell constructions typically result in a first-order phase transition (Christian et al., 2024; Huang & Sourav, 2025; Counsell et al., 2025). However, recent theoretical studies suggest that a continuous transition, or crossover, could be a more accurate description in the QCD framework (Schäfer & Wilczek, 1999; Hirono & Tanizaki, 2019; Fujimoto et al., 2020, 2025). Percolation theory, in particular, supports the idea that the transition from hadronic to QM need not be abrupt. Instead, it could occur as a smooth crossover, consistent with lattice QCD results at high temperatures. The geometric and statistical principles of percolation theory provide valuable insight into quark-hadron continuity, suggesting that for baryon densities in the range $2n_{0}<n_{B}<(4-7)\leavevmode\nobreak\ n_{0}$, the composition of matter should change smoothly from hadronic to QM (Fukushima et al., 2020; Baym et al., 2018).
@@
Recent observations of NSs show similar radii for stars with masses of $2.1$ and $1.4$ solar masses. This finding challenges the onset of a sharp first-order phase transition in the density range slightly above saturation, extending up to the baryon densities found in the cores of two-solar-mass NSs ($n_{B}\approx 4-7\leavevmode\nobreak\ n_{0}$) (Kojo & Suenaga, 2022). This observation supports the hypothesis of hadron-quark continuity, where a weak first-order phase transition may still exist but is not dominant, allowing for a more accurate description of the QCD matter EOS inside NSs (Kojo, 2021).
@@
It is also worth noting that Hensh et al. (2024) recently analyzed both crossover and strong first-order phase transitions in GW signals. They found that the post-merger main frequency $f_{2}$ is lower than predicted by hadronic models with the same tidal deformability ($\Lambda$). They concluded that the detection of both inspiral and post-merger signals could provide evidence for the presence of a crossover phase transition (Hensh et al., 2024).
In light of the aforementioned points, we adopt the Replacement Interpolation Construction (RIC) method (Ayriyan et al., 2018; Ayriyan & Grigorian, 2018; Shahrbaf et al., 2020b) to model the phase transition. This approach uses a second-order polynomial interpolation between the hadronic and quark phases, ensuring a smooth crossover. It assumes a mixed-phase construction, in which both hadronic and QM coexist in the interpolated region. To ensure thermodynamic consistency, the interpolation is performed at the level of the pressure as a function of baryon chemical potential,$P(\mu_{B})$, with all other thermodynamic quantities obtained by differentiating $P(\mu_{B})$ with respect to $\mu_{B}$. The interpolation is carried out on the total thermodynamic quantities, including contributions from both baryonic-QM and the leptonic sector. This ensures a smooth crossover in the full thermodynamic space, consistent with charge neutrality and beta equilibrium.
The hadronic EOS is considered appropriate at lower densities, particularly in the vicinity of nuclear saturation density, where the BM does not reveal its mixed structure. We establish the upper limit of applicability for our hadronic EOS as $n_{H}(\mu_{B})$. In regions of relatively high densities, where quarks exist independently of specific baryons, the deconfined QM EOS is invoked. The lower threshold of validity for the color superconducting QM EOS is defined as $n_{Q}(\mu_{B})$. In the intermediate density range, denoted by $n_{H}(\mu_{B})<n(\mu_{B})<n_{Q}(\mu_{B})$, neither the hadronic EOS nor the QM EOS is suitable.
@@
The determination of surface tension at the interface between the hadronic phase and quark phase in strongly interacting matter remains a topic of ongoing investigations. An infinite surface tension implies the applicability of a Maxwell construction, whereas a vanishing surface tension necessitates the use of a Glendenning construction. In scenarios where the surface tension is variable, a crossover construction approach can be considered.
@@
The EOS for the hadronic and QM phases are expressed in terms of the pressure versus chemical potential, denoted as $P_{H}(\mu_{B})$ and $P_{Q}(\mu_{B})$, respectively, at $T=0$, which is relevant for NSs. Under these conditions, the effective crossover EOS, $P_{M}(\mu_{B})$, can be conveniently represented as an interpolation function between the two phases. This interpolation requires that the pressure function connects smoothly the values characteristic of hadronic and quark phases at their lower and upper limits. Moreover, it must satisfy fundamental thermodynamic constraints, including a positive slope of density versus chemical potential, i.e., $\frac{\partial n_{M}}{\partial\mu_{M}}=\frac{\partial^{2}P_{M}}{\partial\mu_{M}^{2}}>0$. Additionally, the causality condition, which ensures that the adiabatic speed of sound at zero frequency, $c_{s}^{2}=\frac{\partial P}{\partial\epsilon}$, does not exceed the speed of light, must be obeyed too. One effective and practical approach for achieving this interpolation is through the utilization of a polynomial function, which smoothly connects the pressure curves of the two phases.
The critical chemical potential, denoted as $\mu_{c}$, defines the point at which two phases achieve both mechanical and chemical equilibrium. It is determined through the Maxwell point, encompassing both the physical transition and a non-physical transition referred to as reconfinement, expressed as:
@@
To model the pressure of the crossover phase, we employ a polynomial ansatz:
$P_{M}(\mu_{B})=\sum_{q=1}^{N}\alpha_{q}(\mu_{B}-\mu_{c})^{q}+(1+\Delta_{P})P_{c}.$ (9)
Here, $\Delta_{P}$ represents a free parameter that characterizes the pressure of the cross-over phase at $\mu_{c}$:
$P_{M}(\mu_{c})=P_{c}+\Delta_{P}=P_{M},\text{ where }\Delta_{P}=\frac{\Delta P}{P_{c}}.$ (10)
The value of $\Delta_{P}$ is intricately related to the surface tension existing at the interface between the two phases (Maslov et al., 2019). Typically, within the framework of Eq. (9), it is common to consider $N=2$.
@@
$P_{H}(\mu_{H})$ $=$ $P_{M}(\mu_{H})\leavevmode\nobreak\ ,$ (11)
$P_{Q}(\mu_{Q})$ $=$ $P_{M}(\mu_{Q})\leavevmode\nobreak\ ,$ (12)
$\frac{\partial}{\partial\mu_{B}}P_{H}(\mu_{H})$ $=$ $\frac{\partial}{\partial\mu_{B}}P_{M}(\mu_{H})\leavevmode\nobreak\ ,$ (13)
$\frac{\partial}{\partial\mu_{B}}P_{Q}(\mu_{Q})$ $=$ $\frac{\partial}{\partial\mu_{B}}P_{M}(\mu_{Q})\leavevmode\nobreak\ .$ (14)
By considering the relationship $n(\mu_{B})=\frac{\partial P(\mu_{B})}{\partial\mu_{B}}$, we undertake a numerical solution of the continuity equations, i.e., (11-14) for baryon density. This approach enables us to deduce the unknown variables in the aforementioned equations.
It is notable to emphasize that for the employed EOSs, the first crossing between the hadronic phase and the deconfined QM phase takes place at extremely low densities. Consequently, we employ a cross-over construction for the second crossing corresponding to the reconfinement transition from the quark phase to the hadronic phase at higher densities, wherein $\Delta_{P}<0$. This approach allows the hadronic EOS to remain applicable at lower densities, while the QM EOS remains valid at higher densities.
3 EOS Characterization Across Phases and Thermodynamic Behavior
We have determined specific values for the slope parameter, denoted as $x_{\text{S}}$, which corresponds to the coupling strength, and for the vacuum mass of S, $m_{\text{S}}$, in Eq (3), by ensuring they satisfy the constraints arising from the interplay between these two parameters. The resulting constraint region, illustrating $x_{\text{S}}$ as a function of $m_{\text{S}}$, is presented in Figure 8 of our previous study (Shahrbaf et al., 2022).

### v3

@@ 1 Introduction
In our prior work, Shahrbaf et al. (2022), we established the lower-mass limit for S, ensuring that the possible emergence of S as a DM candidate remains consistent with observational constraints derived from NSs. Our findings disclosed the feasibility of S appearance in the hadronic phase of NS, demonstrating that a minimum mass of $m_{\texttt{S}}=1885$ MeV fulfills the necessary conditions that S does not appear at sub-saturation densities. In the present work, we adopt the same approach to ascertain the upper limit for the mass of S, assuming that the interaction strength with BM is fixed at its lowest value. This provides a reliable assumption about the DM-ordinary matter interaction.
@@
Furthermore, we performed a Bayesian analysis (BA) to systematically determine the posterior probability distribution of the allowed mass range of S, accounting for both observational uncertainties and model assumptions. By applying multiple constraints, the Bayesian framework provides more precise limits on the allowed bosonic DM parameter space within the context of NS physics. For our analysis, we considered fundamental constraints on NSs, mass, radius, and tidal deformability along with two extreme cases: HESS J1731-347 (Doroshenko et al., 2022), a light and ultracompact NS, and PSR J0952–0607 (Romani et al., 2022), which is a black widow (BW) pulsar and the most massive confirmed NS. The constraints are first applied independently to infer the respective posterior distributions of the S mass. Subsequently, these are combined to yield the final posterior distribution.
@@
The paper is organized as follows: In Section 2, we introduce our model to describe the hadronic phase with hyperons and DM as well as the QM phase. Our approach for combining these two phases and constructing a crossover phase transition is also discussed in this section. Our findings and in-depth discussions concerning the thermodynamic properties of the resulting hybrid stars are presented in Section 3. In Section 4, we investigate the impact of NS constraints to find the upper limit of the mass of DM particles. Our BA approach based on the NS observations is presented in Section 5. Finally, we summarize our work and draw conclusions in Section 6.
@@ 2.1 Hadronic matter equation of state including hyperons and dark matter
In the DD2Y-T model, all particles are treated as quasi-particles with effective mass and effective chemical potential. The interactions in this model exclusively entail meson exchange interactions, and the couplings to $\sigma$, $\omega$, $\rho$, and $\phi$ mesons depend on the density of the medium. The density-dependent couplings are accurately adjusted to faithfully reproduce the characteristics of atomic nuclei, as established by earlier works (Typel and Wolter, 1999; Typel, 2005; Typel et al., 2010).
The DD2Y-T model has been employed to assess the properties of NSs in both high-temperature (or ”hot”) and low-temperature (or ”cold”) hyperonic configurations (Stone et al., 2021). Notably, it is essential to highlight that this model successfully addresses the ’hyperon puzzle’, a longstanding problem that has been extensively deliberated in the literature (Sedrakian et al., 2020; Choi et al., 2023; Tolos, 2025; Fujimoto et al., 2024; Tong et al., 2025). It is considered one of the most challenging issues within the domain of hypernuclear physics, and various works have tried to address the solutions to this puzzle (Maslov et al., 2015; Shahrbaf and Moshfegh, 2019; Shahrbaf et al., 2020a, b; Li and Sedrakian, 2023; Tsiopelas et al., 2024; Jangal et al., 2025).
@@
As established in Shahrbaf et al. (2022), we have demonstrated that a constant mass assumption for the S particle is incompatible with the structural stability of NS. Consequently, we introduced a linear density-dependent mass shift for the S particle, with a slope parameter harmonized with the corresponding mass shifts observed in other baryons. Detailed constraints regarding the lower threshold of the mass and the slope parameter have been thoughtfully presented and illustrated in Figure 8 of Shahrbaf et al. (2022). In our approach,
$S_{S}=-\Delta m_{S}$ (1)
@@
$\Delta m_{S}=m_{S}x_{S}\frac{n_{B}}{n_{0}},$ (3)
where $n_{B}$ is the baryonic density, $x_{S}$ is a slope parameter, and $n_{0}=0.15$ fm-3 is the nuclear saturation density. For the S particle, we adopted the minimal value of the slope parameter. This choice, while simplified, produces a significant mass shift at high densities of NSs and ensures that the S remains a deeply bound but weakly interacting state. Such a treatment is consistent with the expectation that a viable DM candidate should interact only weakly with BM, and it serves to mimic the possible interactions of the S with the medium without introducing additional free parameters. A finite density of S exclusively emerges at zero temperature when the condition $m^{*}_{S}=\mu^{*}_{S}$ is met, leading to the occurrence of Bose-Einstein condensation. The scalar density for S is then equal to the vector density and reads
@@
where $\mu_{S}^{\ast}=\mu_{S}-V_{S}$ is the effective chemical potential and $\mu_{S}=2\mu_{B}$ is the chemical potential of the S particle. $V_{S}$ is the vector potential for the S particle and reads
@@ 2.3 Hybrid equation of state
In the context of constructing a phase transition to QM, traditional Maxwell constructions typically result in a first-order phase transition (Christian et al., 2024; Huang and Sourav, 2025; Counsell et al., 2025). However, recent theoretical studies suggest that a continuous transition, or crossover, could be a more accurate description in the QCD framework (Schäfer and Wilczek, 1999; Hirono and Tanizaki, 2019; Fujimoto et al., 2020, 2025). Percolation theory, in particular, supports the idea that the transition from hadronic to QM need not be abrupt. Instead, it could occur as a smooth crossover, consistent with lattice QCD results at high temperatures. The geometric and statistical principles of percolation theory provide valuable insight into quark-hadron continuity, suggesting that for baryon densities in the range $2n_{0}<n_{B}<(4-7)\penalty 10000\ n_{0}$, the composition of matter should change smoothly from hadronic to QM (Fukushima et al., 2020; Baym et al., 2018).
@@
Recent observations of NSs show similar radii for stars with masses of $2.1$ and $1.4$ solar masses. This finding challenges the onset of a sharp first-order phase transition in the density range slightly above saturation, extending up to the baryon densities found in the cores of two-solar-mass NSs ($n_{B}\approx 4-7\penalty 10000\ n_{0}$) (Kojo and Suenaga, 2022). This observation supports the hypothesis of hadron-quark continuity, where a weak first-order phase transition may still exist but is not dominant, allowing for a more accurate description of the QCD matter EOS inside NSs (Kojo, 2021).
@@
It is also worth noting that Hensh et al. (2024) recently analyzed both crossover and strong first-order phase transitions in GW signals. They found that the post-merger main frequency $f_{2}$ is lower than predicted by hadronic models with the same tidal deformability ($\Lambda$). They concluded that the detection of both inspiral and post-merger signals could provide evidence of the presence of a crossover phase transition (Hensh et al., 2024).
@@
Moreover, the determination of surface tension at the interface between the hadronic phase and quark phase in strongly interacting matter remains a topic of ongoing investigations. An infinite surface tension implies the applicability of a Maxwell construction, whereas a vanishing surface tension necessitates the use of a Glendenning construction. In scenarios where the surface tension is variable, a crossover construction approach can be considered.
In light of the aforementioned points, we adopt the replacement interpolation construction (RIC) method (Ayriyan et al., 2021, 2018; Ayriyan and Grigorian, 2018; Shahrbaf et al., 2020b) to model the phase transition. Importantly, two distinct interpolation philosophies exist in the literature. In the earlier works of Ayriyan et al. (2018); Ayriyan and Grigorian (2018), the interpolated EOS was considered as a thermodynamically meaningful mixed phase, in which the $\Delta_{P}$ parameter reflects the physical surface tension at the hadron–quark interface. In contrast, in the later formulation of Ayriyan et al. (2021), the interpolation is introduced purely as a mathematical bridge between two trustworthy regions, namely the hadronic EOS at $\mu_{B}<\mu_{H}$ and the quark EOS at $\mu_{B}>\mu_{Q}$. In this latter interpretation, which we explicitly follow in the present work, the interpolated EOS does not represent a thermodynamically favored state. Instead, it is a pragmatic construction enabling us to explore the astrophysical consequences of a smooth crossover rather than a sharp first-order transition. It is worth mentioning that both the baryonic-QM and the leptonic sectors are taken into account in the interpolation procedure.
The hadronic EOS is considered appropriate at lower densities, particularly in the vicinity of nuclear saturation density, where the BM does not reveal its mixed structure. We establish the upper limit of applicability for our hadronic EOS as $n_{H}(\mu_{B})$. In regions of relatively high densities, where quarks exist independently of specific baryons, the deconfined QM EOS is invoked. The lower threshold of validity for the color superconducting QM EOS is defined as $n_{Q}(\mu_{B})$. In the intermediate density range, denoted by $n_{H}(\mu_{B})<n(\mu_{B})<n_{Q}(\mu_{B})$, neither the hadronic EOS nor the QM EOS is suitable, and the interpolated EOS is applied as a mathematical connection between them.
@@
The EOS for the hadronic and QM phases are expressed in terms of the pressure versus chemical potential, denoted as $P_{H}(\mu_{B})$ and $P_{Q}(\mu_{B})$, respectively, at $T=0$, which is relevant for NSs. Under these conditions, the effective crossover EOS, $P_{M}(\mu_{B})$, can be conveniently represented as an interpolation function between the two phases. This interpolation requires that the pressure function connects smoothly the values characteristic of hadronic and quark phases at their lower and upper limits. Moreover, it must satisfy fundamental thermodynamic constraints, including a positive slope of density versus chemical potential, i.e., $\frac{\partial n_{M}}{\partial\mu_{B}}=\frac{\partial^{2}P_{M}}{\partial\mu_{B}^{2}}>0$. Additionally, the causality condition, which ensures that the adiabatic speed of sound at zero frequency, $c_{s}^{2}=\frac{\partial P}{\partial\epsilon}$, does not exceed the speed of light, must be obeyed too. One effective and practical approach for achieving this interpolation is through the utilization of a polynomial function, which smoothly connects the pressure curves of the two phases.
The critical chemical potential, denoted as $\mu_{c}$, is defined as the mathematical crossing point between the hadronic and quark EOSs. While in the Maxwell construction, this point has a thermodynamic meaning, in our RIC implementation, it serves only as a reference point for the polynomial interpolation. This crossing is expressed as
@@
To model the pressure of the crossover phase, we employed a polynomial ansatz:
$P_{M}(\mu_{B})=\sum_{q=1}^{N}\alpha_{q}(\mu_{B}-\mu_{c})^{q}+(1+\Delta_{P})P_{c},$ (9)
with the free parameter, $\Delta_{P}$, quantifying the offset of the interpolated curve at $\mu_{c}$:
$P_{M}(\mu_{c})=P_{c}+\Delta P=P_{M},\text{ where }\Delta_{P}=\frac{\Delta P}{P_{c}}.$ (10)
In this work, we fixed $\Delta_{P}=-0.02$ for simplicity. Given the choice of well-reliable EOSs for both the hadronic and quark phases and assuming the existence of a phase transition, but taking into account that these EOSs by themselves do not provide a reasonable Maxwell-type phase transition, the RIC method with negative $\Delta_{P}$ could be employed. Typically, within the framework of Eq. (9), it is common to consider $N=2$.
@@
$P_{H}(\mu_{H})$ $=$ $P_{M}(\mu_{H})\penalty 10000\ ,$ (11)
$P_{Q}(\mu_{Q})$ $=$ $P_{M}(\mu_{Q})\penalty 10000\ ,$ (12)
$\frac{\partial}{\partial\mu_{B}}P_{H}(\mu_{H})$ $=$ $\frac{\partial}{\partial\mu_{B}}P_{M}(\mu_{H})\penalty 10000\ ,$ (13)
$\frac{\partial}{\partial\mu_{B}}P_{Q}(\mu_{Q})$ $=$ $\frac{\partial}{\partial\mu_{B}}P_{M}(\mu_{Q})\penalty 10000\ .$ (14)
By considering the relationship $n(\mu_{B})=\frac{\partial P(\mu_{B})}{\partial\mu_{B}}$, we undertake a numerical solution of the continuity equations, i.e., Eqs. (11)–(14), for baryon density. This approach enables us to deduce the unknown variables ($\alpha_{1}$, $\alpha_{2}$, $\mu_{H}$, $\mu_{Q}$) in the aforementioned equations.
Finally, we emphasize that for the EOSs employed here, the first crossing between the hadronic and quark curves takes place at very low densities. This crossing is discarded in our construction. Instead, the second crossing, corresponding to the so-called reconfinement transition at higher densities, is adopted as the boundary for the RIC interpolation. This procedure ensures that the hadronic EOS remains applicable at lower densities, while the QM EOS governs the high-density regime, and the interpolated EOS provides a smooth mathematical connection between them.
2.4 Parameter choices and ranges
For completeness, we outline here the chosen parameter ranges and fixed inputs. These serve as the basis for constructing the hadronic, quark, and hybrid phases of our model.
We scanned the S mass over the interval $m_{S}=1890$$2054\penalty 10000\ \mathrm{MeV}$. The mass-shift slope is fixed at the minimal value $x_{S}=0.03$. Quark matter is modeled using two fixed nlNJL parametrizations: the first with $(\eta_{D},\eta_{V})=(0.70,0.10)$ and the second with $(0.70,0.11)$. For both parametrizations, the same squared speed of sound, $c_{s}^{2}=0.44$, is obtained through the established mapping between the nlNJL and CSS parametrizations, i.e., Eq. (C4) in Ref. Shahrbaf et al. (2023). The hadron-quark crossover is constructed using the quadratic RIC; for each combination of a hadronic EOS (with a specific S-particle mass) and a quark EOS set, we determine numerically the matching chemical potentials $(\mu_{H},\mu_{Q})$ and ($\alpha_{1},\alpha_{2}$) in Eq. (9) which control the interpolation shape. The pressure offset is fixed to $\Delta_{P}=-0.02$ and no additional free parameters are introduced. Therefore, the four unknowns $(\alpha_{1},\alpha_{2},\mu_{H},\mu_{Q})$ are determined by solving the four continuity conditions Eqs. (11)-(14). For readability, curves in multi-$m_{S}$ figures are labeled in increasing $m_{S}$, and some plots show subsets of the range.
3 Equation of state characterization across phases and thermodynamic behavior
We have determined specific values for the slope parameter, denoted as $x_{\text{S}}$, which corresponds to the coupling strength, and for the vacuum mass of S, $m_{\text{S}}$, in Eq. (3), by ensuring they satisfy the constraints arising from the interplay between these two parameters. The resulting constraint region, illustrating $x_{\text{S}}$ as a function of $m_{\text{S}}$, is presented in Figure 8 of our previous study (Shahrbaf et al., 2022).